OpenStudy (anonymous):

Need someone to help me with 5 chemistry questions

OpenStudy (anonymous):

would love to help

OpenStudy (aaronq):

Look at their coefficients \(\sf -\dfrac{1}{2}\dfrac{\Delta [SO_3]}{\Delta t}=\dfrac{1}{2}\dfrac{\Delta [SO_2]}{\Delta t}=\dfrac{\Delta [O_2]}{\Delta t}\)

OpenStudy (anonymous):

What should I look for?

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

if their coefficients are the same, then their rate is the same.

OpenStudy (anonymous):

So it's either A or B?

OpenStudy (aaronq):

yep

OpenStudy (anonymous):

I think it's A

OpenStudy (aaronq):

look at their coefficients. I can't really tell you more without giving you the answer.

OpenStudy (anonymous):

Isn't the coefficient 2?

OpenStudy (aaronq):

for \(SO_3\)? yes

OpenStudy (anonymous):

So it's A

OpenStudy (aaronq):

you're saying that the stoichiometric coefficient of \(O_2\) is also 2?

OpenStudy (anonymous):

Oh, I don't know

OpenStudy (aaronq):

Just look at the equation again.

OpenStudy (anonymous):

I don't understand how to figure that out

OpenStudy (aaronq):

you have \(\sf \color{red}2~SO_3\rightleftharpoons \color{red}2~SO_2+ \color{red}1~O_2\) the numbers in red are the stoichiometric coefficients